OpenStudy (anonymous):

|x+y|= |x|+|y| if x and y are both positive? Both negative? What happens if x is negative and y is positive explain.?

OpenStudy (perl):

solve it by cases

OpenStudy (perl):

|x+y|= |x|+|y| Case 1: x>= 0 , y > = 0 x+ y > = 0 x + y = x + y, which is true for all x y, so quadrant 1 is completely shaded